OpenStudy (anonymous):

The volume ( v) of a gas in a container at a constant temperature varies inversely as the pressure (p). If the volume of a gas is 24 cubic centimeters when the pressure is 16 pounds, what is the volume under a pressure of 30 pounds? 45 cubic cm 20 cubic cm 12.8 cubic cm

OpenStudy (anonymous):

@ribhu

OpenStudy (jack1):

not 100% here, but if it varies inversely, does that mean v = k/p... so v * p = k (where k is a constant)

OpenStudy (jack1):

i think that works, so if the volume goes up, the pressure would decrease, yeah?

OpenStudy (anonymous):

No idea.

OpenStudy (jack1):

ok, welll that works here... so using v = k/p (v is volume, p is pressure and k is a constant) find k if v = 24 and p = 16

OpenStudy (anonymous):

I really don't know how to do these types of problems :/

OpenStudy (jack1):

thats why i'm teaching you... if you can do multiplication and division, you can solve these type of problems.

OpenStudy (jack1):

\(\Large V = \frac kP\) if v = 24 and p = 16... what does k = ?

OpenStudy (anonymous):

384

OpenStudy (jack1):

yep, so that's the constant of this container of gas now k stays = 384 and the new pressure (p) = 30, solve volume (V) \(\Large V = \frac kP\)

OpenStudy (anonymous):

64/5?

OpenStudy (jack1):

yep, done ;) also 64/5 in decimals is 12.8

OpenStudy (anonymous):

Ahhh thank you!

OpenStudy (jack1):

sçool

OpenStudy (jack1):

*s'cool